OpenStudy (imnotgoodatmath):

Which property of equality can be used to solve the equation N + 9 = 3 ? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

keep telling yourself that and you never will be good at math! how would you solve this? what would you do?

OpenStudy (imnotgoodatmath):

first, would you subtract?

OpenStudy (anonymous):

yes, great! so what property do think is appropriate?

OpenStudy (imnotgoodatmath):

ok, i found it

OpenStudy (imnotgoodatmath):

1. Reflexive property 2. 3x=21, x=7 3. t/3 * 3 = -3 * 3, t=-9 4. n = -1

OpenStudy (imnotgoodatmath):

is that right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is this a different problem?